OpenStudy (anonymous):

Solve the triangle. A = 21°, C = 105°, c = 5

OpenStudy (jdoe0001):

so.... what are we solving for?

OpenStudy (anonymous):

the other components. Can you check my solution? B = 54°, a ≈ 1.9, b ≈ 4.2

OpenStudy (scorcher219396):

You can easily solve for B, first 180=a+b+c so 180-105-21=54 Then you can use law of sines for this one \[\frac{ sinA }{ a }=\frac{ sinB }{ b }\frac{ sinC }{ c }\] Then set it up with the information you know \[\frac{ \sin21^{o} }{ a }=\frac{ sinB }{ b } = \frac{ \sin105^{o} }{ 5 }\] You can basically ignore the middle part about sinB and b for now, and take the a and c parts and cross multiply, solve for a \[5\sin21^{o}=asin105^{o}\] Divide both sides by sin105, and a= 1.86 (MAKE SURE YOURE IN DEGREES NOT RADIANS) Then do the same for b with the cross multiplying \[5\sin54^{o}=bsin105\] Divide by sin105, b=4.19 Then its always good to make sure your biggest angle and biggest side line up so that you know you didn't do something stupid

OpenStudy (anonymous):

So, was I correct?

OpenStudy (scorcher219396):

Haha yes, you sent your answers before i got a chance to finish. Good work! :)

OpenStudy (anonymous):

Can you help me with something else?

OpenStudy (jdoe0001):

you can always post anew, thus we can revise each other